OpenStudy (anonymous):

if f(x)=4-3x. what is equal to f'(-1). an explanation would be greatly appreciated

OpenStudy (anonymous):

First you find the general derivative, which will be f'(x)=-3

OpenStudy (anonymous):

okay i got that

OpenStudy (anonymous):

So, f'(-1)=-3

OpenStudy (anonymous):

so for any value it will be -3?

OpenStudy (anonymous):

F(-1) means that if there is an X in the equation, it will be replaced by -1. Does that make sense?

OpenStudy (anonymous):

f(x)=4-3x f'(x)=-3 There are no X's so yes, it will always be -3

OpenStudy (anonymous):

kinda so all values of x would equal -3?

OpenStudy (anonymous):

No. The equation would always be -3. If the equation was f'(-1)=-3x+x then the value would be f(-1)=2. does that make sense. you just substitute whatever value that's in the parentheses into an x IF, AND ONLY IF, there is an X in the equation. In your equation's case, you don't have an X so that will always be the value of that equation. I hope that made more sense :D

OpenStudy (anonymous):

thanks it did!